OpenStudy (anonymous):

ill give you a medal..Name the property shown (4-3x)*5x^2=4*5x^2-3*5x^2) a) inverse property of multiplication b) zero property of multiplication c) identity property of multiplication d) distributive property

OpenStudy (mathstudent55):

I think you meant this: (4-3x)*5x^2=4*5x^2-3x*5x^2 (you left out the x at the 3x)

OpenStudy (anonymous):

oh yess srry

OpenStudy (mathstudent55):

inverse property: a * 1/a = 1 (for a =/= 0) zero property: a * 0 = 0 identity property: a * 1 = a distributive property: a(b + c) = ab + ac Now that you see what the properties are, can you figure it out yourself?

OpenStudy (anonymous):

would it be distributive